OpenStudy (anonymous):

how do you find an for the given geometric sequence! a3=9 r=-3 n=7

OpenStudy (imstuck):

You first need to find a1 using the info they gave you so you can write your equation correctly. Solve for a1 by replacing an with 9, and r with -3. You find a1 = 1.\[a _{n} =a _{1}\times r ^{n-1}\]\[9=a _{1}\times (-3)^{3-1}\]\[9=a _{1}\times (-3)^{2}\]\[9=a _{1}\times9\]\[9=9a _{1}\]\[a _{1}=1\]Now write your sequence using the info for a7. Isn't that what you want if n = 7?

OpenStudy (imstuck):

\[a _{7}=1\times (-3)^{7-1}\]\[a _{7}=1\times (-3)^{6}\]\[a _{7}=1\times 729\]so\[a _{7}=729\]The seventh term in your sequence is 729

OpenStudy (anonymous):

Ohhh!! okay okay. Thank you so much!!

OpenStudy (texaschic101):

an = a1 * r ^(n-1) n = the term you want to find = 7 a1 = first term = 1 r = common ratio = -3 now just sub and you should get the answer that IMStuck got

OpenStudy (imstuck):

TY for the medal!